### v1

# arXiv:2603.02009v1[math.AP] 02 Mar 2026

Abstract.
This paper investigates the critical quintic wave equation in a 3D bounded domain subject to locally distributed Kelvin-Voigt damping. The study tackles two major mathematical challenges: the severe loss of derivatives induced by the localized thermo-viscous dissipation and the aggressive nature of the critical nonlinear term. First, we establish a robust well-posedness theory for arbitrarily large initial data. By shifting the analysis to the frequency space via a Littlewood-Paley decomposition , we employ Bernstein’s inequalities to lift the damping term into an $L^{2}$ framework, allowing Strichartz estimates to be applied flawlessly. In the second part, we prove the uniform exponential stabilization of the energy. To overcome the reduction of the residual to the $H^{-2}$ level caused by the Kelvin-Voigt mechanism, we utilize the microlocal defect measure framework. The core of our stabilization proof relies on combining the critical Strichartz regularity $L^{4}_{t}L^{12}_{x}$ with a sharp Unique Continuation Property (UCP) to close the observability argument. Furthermore, we demonstrate that this microlocal mechanism is perfectly compatible with non-invasive damping geometries of arbitrarily small Lebesgue measure, successfully circumventing the geometric obstruction of trapped rays.
Key words and phrases:
Wave equation, Nonlinear Kelvin-Voigt damping, Critical exponent, Strichartz estimates, Parabolic regularity.
2020 Mathematics Subject Classification:
35L05, 35L70, 35B40, 93D15
1. Introduction
1.1. The phenomenon of energy concentration
The concentration phenomena associated with the "quintic wave" (generally described by the Nonlinear Quintic Schrödinger Equation - NLS or the Semilinear Quintonic Wave Equation) refer to situations where the energy or mass of the wave accumulates intensely in a very small region of space (a point or a small neighborhood), often leading to what is called a "blow-up" (singularity in finite time). From a physical point of view, this represents the collapse of a wave packet, while mathematically it corresponds to a concentration of mass (norm) or energy, where the solution profile loses compactness in Sobolev spaces. Mass/Energy Concentration: Mathematically, a sequence of solutions is said to concentrate if the mass (norm) or energy (norm) concentrates at a point when the concentrated mass is often a discrete quantity, equal to the norm of the "ground state". There is a threshold profile, namely, scattering threshold(the center-stable manifold), that separates solutions that scatter (go to zero) from those that concentrate and collapse (blow-up). Concentration is analyzed through scaling transformations (scaling invariant), where the solution is rescaled to maintain a constant energy norm while the wave packet size goes to zero.
The phenomenon of "collapse" or blow-up in quintic wave equations (and other critical nonlinear wave equations) can perfectly occur in bounded domains. In fact, the presence of edges (bounded domains) can facilitate blow-up compared to free space, as confinement prevents energy from dispersing to infinity, forcing it to concentrate. Here is the detailed analysis based on mathematical physics:
1. Collapse in Bounded Domains.
From a physical point of view, in a bounded domain, the initial energy "trapped" in the volume, combined with the non-linear focus (typical of the quintic wave), can overcome the linear mechanisms of dispersion. The wave "collapses" because the energy concentrates at a point (or region) in finite time, resulting in the singularity.
2. Mathematical View and Sobolev Spaces.
Mass/Energy Concentration: Mathematical collapse is the blow-up of the energy norm or norms in Sobolev spaces, such as:
  • As time approaches the collapse time, the solution ceases to be compact, meaning that the gradient of the solution becomes infinite, concentrating at one or more points.
  • Loss of Compactness: In bounded domains, immersion theorems (such as Rellich-Kondrachov) guarantee initial compactness, but the quintic nature of nonlinearity is "critical" in energy space. When the blow-up occurs, this compactness is violated, as the solution profile becomes singular (tends to a Dirac delta-type distribution in the limit).
3. Factors Affecting Collapse:
Boundary: The shape of the domain and boundary conditions (e.g., Dirichlet at the edge) influence the rate of energy concentration, but collapse is an interior concentration phenomenon, i.e., it occurs where nonlinearity is strong.
Damping: If the system includes damping terms (damped wave equations), it is possible for the system to converge to a global attractor, avoiding collapse, depending on the intensity of the damping.
Nonlinear Equations: Collapse is characteristic of focusing equations (where nonlinearity attracts the wave), while in defocusing equations, the wave disperses.
In summary, the quintic wave in bounded domains frequently exhibits collapse (finite-time singularity) where the energy norm in Sobolev space explodes, a phenomenon that can be exacerbated by the inability of energy to escape the domain.
1.2. The drawback of Galerkin’s method
Since we are in a bounded domain and with the spectral basis of the Laplacian, the weak existence and regularity of the approximate problem (via Galerkin) are practically "free" due to compactness and Aubin-Lions. The term $u^{5}\in L^{6/5}$ is perfectly tamed by the Sobolev immersion $H_{0}^{1}\hookrightarrow L^{6}$.
The effective linear term in the interval $I$ for the approximation $m$ is:
$A_{m}(I)=\|S(t-t_{0})(u_{m}(t_{0}),\partial_{t}u_{m}(t_{0}))\|_{L^{5}_{t}L^{10}_{x}(I)}.$
We have, for a fixed $m$, that, by the absolute continuity of the Lebesgue integral, we can shrink the interval $I$ such that $A_{m}(I)\leq\epsilon_{0}$. Thus, we divide $[0,T]$ into $N_{m}$ subintervals.
The major problem is in the transition to the limit $m\rightarrow\infty$: For the final bound $\|u_{m}\|_{L^{5}_{t}L^{10}_{x}([0,T])}\leq N\cdot C(E_{0})^{5}$ not to explode when we take the supremum at $m$, the number of subintervals $N$ must be independent of $m$. This means that we need a uniform $\delta>0$ (independent of $m$) such that, if $|I|<\delta$, then $A_{m}(I)\leq\epsilon_{0}$ for all $m$.
In the subcritical case (for example, cubic power in $3D$), we achieve this uniform $\delta$ using subcritical Sobolev immersions interpolated with the energy $E_{0}$. Energy limits the Strichartz norm uniformly within small intervals (see, for instance, [8]).
In the critical case (quintic power in $3D$, norm $L^{5}L^{10}$), the Strichartz exponent "scales" exactly like the energy. The fact that the energy $E_{0}(u_{m})$ is uniformly bounded does not prevent concentration. As the base $m$ grows, the initial data $(u_{m}(0),\partial_{t}u_{m}(0))$ can form increasingly finer "peaks" (energy concentration).
For a thin peak, the energy $H^{1}$ remains the same (limited by $E_{0}$). But the norm $L^{5}_{t}L^{10}_{x}$ of the linear evolution concentrates in an increasingly smaller time interval around $t_{0}$. Therefore, to keep $A_{m}(I)\leq\epsilon_{0}$, the size of the interval $\delta_{m}$ has to shrink to zero when $m\to\infty$. Consequently, the number of partitions $N_{m}\to\infty$, and the bound $N_{m}C(E_{0})^{5}$ explode. It fails to extract a weak bound in $L^{5}_{t}L^{10}_{x}$ because the Galerkin bound is not uniform. This is precisely why the traditional Galerkin method fails to give uniqueness for the critical case, even in bounded domains.
To salvage this situation, we have two direct solutions:
Output 1: Small Data. If we assume that the initial energy $E_{0}$ is small enough, the global Strichartz estimate gives us:
$A_{m}([0,T])\leq CE_{0}\leq\epsilon_{0}$
uniformly for all $m$. We don’t need to divide the interval. $N=1$. And the usual Bootstrap lemma closes the proof and we obtain the uniform bound $L^{5}L^{10}$ in $m$. Then it’s just a matter of passing the limit and it’s over.
Output 2: If Large Data is needed (we have to change the Galerkin method to the Strichartz norm). If we want uniqueness for large initial data, we prove the existence of the weak solution by Galerkin (as we already know it works). But uniqueness in $L^{5}L^{10}$ does not come from estimating $u_{m}$. Instead, the existence of solutions in this critical space (which guarantees uniqueness) is generally achieved by constructing the local solution via the Picard Fixed-Point Theorem directly in the Strichartz space $C([0,\delta];H_{0}^{1})\cap L^{5}([0,\delta];L^{10})$. The fixed point gives us a local existence time $\delta$ that depends on the profile of the initial data, and not just on $E_{0}$. (To extend this $\delta$ to a global $T$ for large data, then the heavy non-concentration arguments would come into play).
In summary: Weak existence proof using Galerkin and compactness works perfectly. But if the intention is to prove that the Galerkin limit is $L^{5}L^{10}$ to guarantee uniqueness with large data, partitioning fails because $N$ goes to infinity with $m$. If it’s for small data, partitioning is unnecessary and Bootstrap shines.
Remark 1.1.
We would like to take this opportunity to mention that, in the manuscript [9] by the authors of present article, we assumed, without any explicit mention, that Galerkin’s method would work well for small datasets and that, for large datasets, we would need to prove or assume that we have a Shatah and Struwe solution, as done in [11], [20], although we did cite those references. Since our primary focus was on decay, we ended up not mentioning this fact. Therefore, we would like to apologize to the readers for this oversight.
The article by Burq, Lebeau, and Planchon (in short BLP) [4] or Blair, Smith, and Sogge (BSS) [3] does not guarantee the uniform boundedness of Strichartz estimates for the Galerkin method approximations. Let’s explain exactly where the calculation stalls and why the literature avoids Galerkin at this specific step, going straight to the ’wound’ of harmonic analysis: The Spectral Projector Problem in $L^{p}$. In the classical Galerkin method, we project the equation onto a finite-dimensional space using the orthogonal projector $P_{m}$, which is an "abrupt" spectral cut on the first $m$ eigenvectors of the Dirichlet Laplacian. To obtain a Strichartz estimate for the approximate solution $u_{m}$, you apply Duhamel’s formula to the projected equation:
$\partial_{tt}u_{m}-\Delta u_{m}+P_{m}(u_{m}^{5})=0.$
The problem lies in the non-linear term $P_{m}(u_{m}^{5})$. To close the estimate, we use the Strichartz bounds of the linear operator, which the BLP actually proves to hold in bounded domains. For this, we need the projector $P_{m}$ to be uniformly bounded in the spatial $L^{p}$ spaces required by the Strichartz duals (for example, something other than $p=2$).
This brings us to a classic and fundamental theorem of harmonic analysis (which goes back to the Fefferman ball multiplier problem [15]), and was exhaustively studied by Christopher Sogge and collaborators in domains/manifolds): The abrupt spectral projector $P_{m}$ is not uniformly bounded on $L^{p}$ for $p\neq 2$. The operator norm $\|P_{m}\|_{\mathcal{L}(L^{p},L^{p})}$ explodes when $m\to\infty$. Therefore, we cannot quote the projected nonlinearity independently of $m$. The constant of the Strichartz estimate from Galerkin’s method for $u_{m}$ will depend on $m$ and explode in the limit. The famous BLP paper (2008, in JAMS) proves the Strichartz estimates for the linear equation in domains with a boundary (overcoming wave reflection). But to deal with frequency cutoffs, they don’t use abrupt cutoffs like in Galerkin. They use pseudo-differential operators and smooth multipliers (Littlewood-Paley type truncations). The smooth cutoff function "damps" the oscillation and ensures the bound on $L^{p}$. Furthermore, to construct the solution to the critical nonlinear problem, BLP (as well as Shatah-Struwe in $\mathbb{R}^{3}$) does not use the Galerkin approximation. They construct the local solution directly in the Strichartz space $L^{5}_{t}L^{10}_{x}$ using Picard’s Fixed Point Theorem, and then use energy non-concentration techniques (Morawetz type) to extend the solution globally.
1.3. Literature Overview
The stabilization of wave equations with localized dissipation has been extensively studied within the geometric control framework initiated by Bardos, Lebeau, and Rauch [2], where sharp conditions linking observability, control, and decay were established. Subsequent developments clarified the microlocal propagation mechanisms governing energy concentration, particularly near the boundary, through the analysis of generalized bicharacteristics [5, 22, 23, 24, 25, 26, 21]. Particularly those involving inhomogeneous media, variable coefficients, or complex boundary configurations, the classical localized GCC fails due to the existence of trapped rays. These are generalized bicharacteristics that oscillate indefinitely without ever entering a standard localized control region, allowing high-frequency energy to concentrate and consequently preventing uniform exponential decay.
In the presence of Kelvin–Voigt damping, however, the analytical structure changes significantly. The damping operator introduces higher-order spatial effects which interact nontrivially with geometric multipliers, frequently requiring additional regularity assumptions. This phenomenon has been examined in various stabilization settings, including localized and degenerate dissipation mechanisms (see [1]).
From the dispersive viewpoint, Strichartz estimates play a decisive role in handling critical nonlinearities. In domains with boundary, these estimates require refined harmonic analysis tools, as developed by Blair, Smith, and Sogge [3], extending earlier Euclidean results of Keel and Tao [19]. The interaction between boundary geometry and dispersive smoothing was further explored by Burq, Lebeau, and Planchon [4], providing a fundamental framework for energy-critical wave equations.
Microlocal defect measures and H-measures provide the natural language for describing high-frequency energy concentration in such problems. These tools, introduced by Gérard [16] and Tartar [30], have become central in stabilization theory, particularly in configurations where compactness arguments fail or where dissipation degenerates.
The present work combines these perspectives. By integrating harmonic analysis, nonlinear energy methods, and microlocal propagation techniques, we obtain a stability theory capable of handling Kelvin–Voigt mechanisms at the critical level while preserving the geometric intuition underlying classical control results.
1.4. Contributions of the present article
The study of the critical wave equation with localized Kelvin-Voigt damping presents a notorious challenge: balancing the aggressive nature of the quintic nonlinearity in $3$ spatial dimensions with the severe loss of derivatives induced by the viscous term $\text{div}(a(x)\nabla u_{t})$. In the first part of this paper, we explore the classical route. By relying on energy methods and seeking strong solutions, one can seemingly bypass the heavy machinery of dispersive estimates. However, this classical approach exacts a heavy toll: to control the critical nonlinearity and guarantee uniqueness, the initial data must be assumed to be strictly small, and the solutions are forced into the restrictive Shatah-Struwe class, requiring $u\in L^{4}(0,T;L^{12}(\Omega))$.
To overcome this bottleneck and remove the small data restriction, the second part of our work shifts the paradigm from the physical domain to the frequency space, deploying the modern arsenal of Harmonic Analysis. The application of dispersive techniques to boundary value problems possesses a rich and profound history. The seminal works of Smith and Sogge [27] demonstrated how wave dispersion interacts with domains and rough boundaries. Subsequently, the celebrated contributions of the trio Burq, Lebeau, and Planchon (BLP) solidified the use of spectral projectors and Littlewood-Paley theory tailored to Dirichlet boundary conditions, allowing Strichartz estimates to flourish outside of $\mathbb{R}^{n}$. Furthermore, treating the singular damping terms via the Duhamel formulation crucially relies on the foundational machinery of Christ and Kiselev, which elegantly translates homogeneous dispersive bounds into retarded estimates, allowing low-frequency perturbations to be seamlessly absorbed.
The core difficulty of the Kelvin-Voigt term lies in its $H^{-1}$ regularity, which makes it inherently hostile to the classical Strichartz estimates on the right-hand side of the wave equation. Classical Strichartz estimates are calibrated for data and source terms with compatible scaling and regularity. The Kelvin-Voigt operator introduces a derivative loss which is incompatible with a direct application of the dispersive framework. The obstruction is therefore not merely dispersive, but functional-analytic in nature.
By introducing a Littlewood-Paley decomposition, we separate the wave into macroscopic (low-frequency) and microscopic (high-frequency) components. This dual-frequency mechanism is the central stabilizing principle of our nonlinear analysis:
  • The Low-Frequency Regime: At low frequencies, the spectral projection $P_{\leq N}$ combined with Bernstein’s inequalities converts spectral localization into effective derivative gains. In this regime, the projected damping term behaves as a bounded perturbation in $L^{2}$, thereby completely neutralizing the derivative loss and allowing the Strichartz estimates to perform flawlessly without destroying the regularity.
  • The High-Frequency Regime: Conversely, at high frequencies, Bernstein’s inequality cannot be applied without yielding unfavorable positive powers of the frequency threshold $N$. Instead, our strategy exploits the finiteness of the initial energy. Since the high-frequency tail of the initial data converges to zero as $N\to\infty$, its contribution to the critical Strichartz norm can be made arbitrarily small. Thus, the high-frequency regime is controlled not by lifting the derivative, but by the asymptotic smallness of the energy tail.
This spectral strategy yields our second and main result: a robust and complete local well-posedness theory. Through a delicate bootstrap argument—combining the smallness of the high-frequency tail with the smooth behavior of the low-frequency component over a short time—we completely bypass the need for small initial energy. Consequently, we establish the existence and uniqueness of solutions for arbitrarily large initial data, providing a definitive structural answer to the conflict between the critical nonlinear power and the Kelvin-Voigt dissipation.
Having established the global well-posedness framework, the third and final objective of this paper is to prove the uniform exponential stabilization of the system. The presence of the Kelvin-Voigt damping $\text{div}(a(x)\nabla u_{t})$ restricted to a localized region $\omega$ introduces severe obstacles for the asymptotic analysis. Unlike frictional damping, the viscous nature of our dissipation alters the regularity of the residual, reducing the convergence from the $H^{-1}$ to the $H^{-2}$ level. Consequently, classical compactness-uniqueness arguments fail to provide the necessary strong convergence to propagate the damping effects.
To bypass this obstacle, we employ the powerful framework of microlocal defect measures. While the loss of derivatives prevents direct propagation everywhere, the microlocal structure is perfectly preserved outside the damping region. The fundamental breakthrough in our stabilization proof is the application of the sharp Unique Continuation Property (UCP) for the wave equation with critical potentials, as established by Duyckaerts, Zhang, and Zuazua [14]. By exploiting the precise critical Strichartz regularity $u\in L^{4}(0,T;L^{12}(\Omega))$ inherited from our well-posedness theory, we can close the defect-measure contradiction argument directly. This completely eliminates the dependence on the truncation parameters without relying on artificial secondary contradictions, bridging the Harmonic Analysis results from the first part of the paper with the Control Theory goals of the second.
Finally, we demonstrate that our microlocal approach is entirely oblivious to volumetric constraints, relying solely on the dynamical interception of the generalized bicharacteristic flow. This allows our stabilization results to be seamlessly coupled with the highly complex, “non-invasive” damping geometries developed in [6, 7, 8]. In such configurations, the dissipative region can be constructed with an arbitrarily small Lebesgue measure, thus providing a definitive and elegant mathematical answer to the problem of trapped rays in inhomogeneous media.
The remainder of the paper is organized as follows. Section 2 constructs global weak solutions via $C^{1}$ truncation. Section 3 develops the local theory of strong solutions through Galerkin methods and higher-order energy estimates. Section 4 establishes uniqueness and weak-strong uniqueness principles. Section 5 introduces the Littlewood-Paley strategy and the frequency-space stabilization of the nonlinear estimates. Section 6 concludes with the microlocal analysis and the proof of uniform exponential decay.
2. Global Weak Solutions via $C^{1}$ Truncation
We consider the critical quintic wave equation with localized Kelvin-Voigt damping:
$\begin{cases}u_{tt}-\Delta u-\hbox{div}(a(x)\nabla u_{t})+u^{5}=0&\text{in }\Omega\times(0,T),\\ u=0&\text{on }\partial\Omega\times(0,T),\\ u(0)=u_{0},\quad u_{t}(0)=u_{1}&\text{in }\Omega,\end{cases}$ (1)
where $\Omega\subset\mathbb{R}^{3}$ is a bounded domain, and $a\in W^{1,\infty}(\Omega)$ with $a(x)\geq 0$. The initial data satisfy $(u_{0},u_{1})\in H_{0}^{1}(\Omega)\times L^{2}(\Omega)$.
2.1. Step 1: The $C^{1}$ Truncated Problem
To apply the Galerkin method without dealing with critical polynomial growth, we truncate the nonlinearity. For each $k\in\mathbb{N}$, define the function $f_{k}:\mathbb{R}\to\mathbb{R}$ by:
$f_{k}(s)=\begin{cases}s^{5}&\text{if }|s|\leq k,\\ \text{sgn}(s)\big[k^{5}+5k^{4}(|s|-k)\big]&\text{if }|s|>k.\end{cases}$ (2)
Lemma 2.1 (Properties of $f_{k}$).
For each $k\in\mathbb{N}$, the following properties hold:
  • (i)
    $f_{k}\in C^{1}(\mathbb{R})$ and is globally Lipschitz with $|f_{k}^{\prime}(s)|\leq 5k^{4}$.
  • (ii)
    $sf_{k}(s)\geq 0$ for all $s\in\mathbb{R}$.
  • (iii)
    $|f_{k}(s)|\leq|s|^{5}$ for all $s\in\mathbb{R}$.
  • (iv)
    $f_{k}(s)\to s^{5}$ uniformly on any compact set $K\subset\mathbb{R}$ as $k\to\infty$.
Proof.
(i) Clearly $f_{k}$ is continuous. The derivative is $5s^{4}$ for $|s|\leq k$ and $5k^{4}$ for $|s|>k$, which matches continuously at $|s|=k$. Thus $f_{k}\in C^{1}(\mathbb{R})$ and $\sup|f_{k}^{\prime}|=5k^{4}$.
(ii) By definition, $f_{k}(s)$ shares the same sign as $s$.
(iii) For $|s|\leq k$, $|f_{k}(s)|=|s|^{5}$. For $s>k$, let $h(s)=s^{5}-(k^{5}+5k^{4}(s-k))$. Since $h(k)=0$ and $h^{\prime}(s)=5(s^{4}-k^{4})>0$ for $s>k$, we have $h(s)>0\implies f_{k}(s)\leq s^{5}$. By symmetry, $|f_{k}(s)|\leq|s|^{5}$ for all $s$.
(iv) Let $K\subset[-M,M]$. For any $k>M$, $|s|\leq k$ for all $s\in K$, which implies $f_{k}(s)=s^{5}$ exactly on $K$. Thus, $\sup_{s\in K}|f_{k}(s)-s^{5}|=0$ for all $k>M$, yielding uniform convergence. ∎
We define $F_{k}(s)=\int_{0}^{s}f_{k}(r)dr$. Property (ii) ensures $F_{k}(s)\geq 0$, and property (iii) ensures $F_{k}(s)\leq\frac{1}{6}|s|^{6}$.
2.2. Step 2: Galerkin Method for the Truncated Problem
For a fixed $k$, we consider the subcritical problem:
$u_{tt}-\Delta u-\hbox{div}(a(x)\nabla u_{t})+f_{k}(u)=0.$ (3)
Let $V_{m}=\text{span}\{\omega_{1},\dots,\omega_{m}\}$ be the space generated by the first $m$ eigenfunctions of the Dirichlet Laplacian. We seek an approximate solution $u_{m}(t,x)=\sum_{j=1}^{m}g_{jm}(t)\omega_{j}(x)$ satisfying the projected system:
$(u_{m}^{\prime\prime}(t),\omega_{j})_{\Omega}+(\nabla u_{m},\nabla\omega_{j})_{\Omega}+(a(x)\nabla u_{m}^{\prime},\nabla\omega_{j})_{\Omega}+(f_{k}(u_{m}),\omega_{j})_{\Omega}=0,\quad j=1,\dots,m,$ (4)
with initial data $u_{m}(0)=P_{m}u_{0}\to u_{0}$ in $H_{0}^{1}$ and $u_{m}^{\prime}(0)=P_{m}u_{1}\to u_{1}$ in $L^{2}$.
Since $f_{k}$ is globally Lipschitz (Lemma 2.1(i)), the mapping $u_{m}\mapsto P_{m}f_{k}(u_{m})$ is globally Lipschitz continuous in $V_{m}$. By the Picard-Lindelöf theorem, the ODE system (4) admits a unique global strong solution $g_{jm}\in C^{2}([0,T])$.
Testing (4) with $g_{jm}^{\prime}(t)$ and summing over $j$, we obtain the identity:
$\frac{d}{dt}\left[\frac{1}{2}\|u_{m}^{\prime}(t)\|_{2}^{2}+\frac{1}{2}\|\nabla u_{m}(t)\|_{2}^{2}+\int_{\Omega}F_{k}(u_{m}(t))dx\right]+\int_{\Omega}a(x)|\nabla u_{m}^{\prime}(t)|^{2}dx=0.$ (5)
Let $E_{m}(t)$ denote the bracketed term. Since $F_{k}(s)\leq\frac{1}{6}|s|^{6}$, we have $E_{m}(0)\leq\frac{1}{2}\|P_{m}u_{1}\|_{2}^{2}+\frac{1}{2}\|\nabla P_{m}u_{0}\|_{2}^{2}+\frac{1}{6}\|P_{m}u_{0}\|_{6}^{6}\leq C_{0}$, where $C_{0}$ depends only on $\|(u_{0},u_{1})\|_{H_{0}^{1}\times L^{2}}$ and not on $k$ or $m$. Integration over $(0,t)$ yields:
$\|u_{m}^{\prime}(t)\|_{2}^{2}+\|\nabla u_{m}(t)\|_{2}^{2}+\int_{0}^{T}\int_{\Omega}a(x)|\nabla u_{m}^{\prime}|^{2}dxdt\leq C_{0}.$ (6)
Because the bound $C_{0}$ is independent of $m$, we can extract a subsequence such that:
$u_{m}$ $\rightharpoonup u_{k}\text{ weakly-$\ast$ in }L^{\infty}(0,T;H_{0}^{1}(\Omega)),$ (7)
$u_{m}^{\prime}$ $\rightharpoonup u_{k}^{\prime}\text{ weakly-$\ast$ in }L^{\infty}(0,T;L^{2}(\Omega)),$ (8)
$u_{m}$ $\to u_{k}\text{ strongly in }L^{2}(0,T;L^{2}(\Omega))\text{ and a.e. in }\Omega\times(0,T).$ (9)
Since $f_{k}$ is globally Lipschitz, $|f_{k}(u_{m})|\leq 5k^{4}|u_{m}|\in L^{\infty}(0,T;L^{2}(\Omega))$. The strong convergence in $L^{2}$ guarantees that $f_{k}(u_{m})\to f_{k}(u_{k})$ strongly in $L^{2}(\Omega\times(0,T))$. Passing to the limit $m\to\infty$ in (4) is straightforward, yielding a global weak solution $u_{k}$ to the truncated problem (3). By lower semicontinuity, $u_{k}$ satisfies the identical energy bound (6), entirely independent of $k$.
2.3. Step 3: The Critical Limit $k\to\infty$
We now have a sequence of exact weak solutions $u_{k}$ to (3), satisfying:
$\sup_{t\in[0,T]}\left(\|u_{k}^{\prime}(t)\|_{2}^{2}+\|\nabla u_{k}(t)\|_{2}^{2}\right)+\int_{0}^{T}\int_{\Omega}a(x)|\nabla u_{k}^{\prime}|^{2}dxdt\leq C_{0}.$ (10)
By the Aubin-Lions lemma, there exists a subsequence $u_{k}\rightharpoonup u$ weakly-$\ast$ in $L^{\infty}H_{0}^{1}$ and strongly a.e. in $\Omega\times(0,T)$.
The Nonlinearity: By the 3D Sobolev embedding $H_{0}^{1}\hookrightarrow L^{6}$, the uniform bound (10) implies that $\{u_{k}\}$ is uniformly bounded in $L^{\infty}(0,T;L^{6}(\Omega))$. Exploiting Lemma 2.1(iii), we have:
$\int_{\Omega}|f_{k}(u_{k}(t,x))|^{6/5}dx\leq\int_{\Omega}|u_{k}(t,x)|^{6}dx\leq C\|u_{k}(t)\|_{H_{0}^{1}}^{6}\leq C_{1},$ (11)
where $C_{1}$ is independent of $k$ and $t$. Thus, $\{f_{k}(u_{k})\}$ is uniformly bounded in the reflexive Banach space $L^{\infty}(0,T;L^{6/5}(\Omega))$.
By the Banach-Alaoglu theorem, we can extract a subsequence such that $f_{k}(u_{k})\rightharpoonup\chi$ weakly-$\ast$ in $L^{\infty}L^{6/5}$. Since $u_{k}\to u$ a.e. and $f_{k}\to s^{5}$ uniformly on compact sets (Lemma 2.1(iv)), it is easy to see that $f_{k}(u_{k}(t,x))\to u(t,x)^{5}$ a.e. in $\Omega\times(0,T)$.
By the standard Lions’ Lemma (Vitali’s convergence theorem for weakly convergent sequences), the weak limit coincides with the pointwise limit:
$f_{k}(u_{k})\rightharpoonup u^{5}\text{ weakly-$\ast$ in }L^{\infty}(0,T;L^{6/5}(\Omega)).$ (12)
Passing to the limit $k\to\infty$ in the weak formulation of (3) is now immediate, since test functions $\phi\in L^{1}(0,T;L^{6}(\Omega))$ exactly dualize the convergence of the nonlinearity. Thus, $u$ is a global weak solution of the critical problem (1).
Theorem 2.2 (Existence of Global Weak Solutions).
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with smooth boundary, and assume that $a\in W^{1,\infty}(\Omega)$ is a non-negative function. For any given initial data $(u_{0},u_{1})\in H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ and any $T>0$, there exists at least one global weak solution $u$ to the critical Kelvin-Voigt damped wave equation (1).
Furthermore, the solution belongs to the class:
$u$ $\in L^{\infty}(0,T;H_{0}^{1}(\Omega)),$
$u_{t}$ $\in L^{\infty}(0,T;L^{2}(\Omega)),$
$\sqrt{a(x)}\nabla u_{t}$ $\in L^{2}(0,T;L^{2}(\Omega)),$
and satisfies the following energy inequality for almost every $t\in(0,T)$:
$E_{u}(t)+\int_{0}^{t}\int_{\Omega}a(x)|\nabla u_{t}(s,x)|^{2}\,dxds\leq E_{u}(0),$ (13)
where the energy is defined as $E_{u}(t)=\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{1}{2}\|\nabla u(t)\|_{2}^{2}+\frac{1}{6}\|u(t)\|_{6}^{6}$.
3. Local Well-posedness of Strong Solutions via Galerkin Method and Small Data
In this section, we establish the local existence of strong solutions for the quintic wave equation with localized Kelvin-Voigt damping. Instead of relying on Strichartz estimates—which present severe microlocal obstructions when commuting with a continuous cut-off function over the localized parabolic damping—we explore the intrinsic degenerate parabolic structure of the system for regular initial data.Let us consider the standard Galerkin approximation for our problem. Let $(w_{j})_{j\in\mathbb{N}}$ be the spectral basis of the Dirichlet Laplacian in $\Omega$, and let $V_{m}=\text{span}\{w_{1},w_{2},\dots,w_{m}\}$. We look for approximate solutions of the form $u_{m}(t,x)=\sum_{j=1}^{m}g_{jm}(t)w_{j}(x)$ solving the projected system:
$(u_{m}^{\prime\prime}(t),w)_{\Omega}+(\nabla u_{m}(t),\nabla w)_{\Omega}+(a(x)\nabla u_{m}^{\prime}(t),\nabla w)_{\Omega}+(u_{m}^{5}(t),w)_{\Omega}=0,\quad\forall w\in V_{m},$
with initial data $u_{m}(0)=P_{m}u_{0}\to u_{0}$ in $H^{2}(\Omega)\cap H_{0}^{1}$ and $u_{m}^{\prime}(0)=P_{m}u_{1}\to u_{1}$ in $H_{0}^{1}(\Omega)$. Before proceeding to the strong energy estimates, we recall that the standard weak dissipative energy:
$E_{m}(t):=\frac{1}{2}\|u_{m}^{\prime}(t)\|_{L^{2}}^{2}+\frac{1}{2}\|\nabla u_{m}(t)\|_{L^{2}}^{2}+\frac{1}{6}\|u_{m}(t)\|_{L^{6}}^{6}$
is non-increasing, satisfying $E_{m}(t)\leq E_{m}(0)\leq E_{0}$ for all $t\geq 0$. In particular, this provides the uniform bound $\|\nabla u_{m}(t)\|_{L^{2}}\leq\sqrt{2E_{0}}$, which will be crucial to control the nonlinear term later.
Defining the Strong Energy.
Since $u_{m}\in V_{m}$ and the basis is given by the eigenfunctions of the Laplacian, we are allowed to take $w=-\Delta u_{m}^{\prime}\in V_{m}$ as a test function in our approximate equation. This choice allows us to capture the regularizing effect of the Kelvin-Voigt damping. We define the strong energy functional as:
$Y_{m}(t):=\frac{1}{2}\|\nabla u_{m}^{\prime}(t)\|_{L^{2}}^{2}+\frac{1}{2}\|\Delta u_{m}(t)\|_{L^{2}}^{2}.$
Substituting $w=-\Delta u_{m}^{\prime}$ into the projected equation and integrating by parts over $\Omega$, we obtain:
$\frac{1}{2}\frac{d}{dt}\left(\|\nabla u_{m}^{\prime}\|_{L^{2}}^{2}+\|\Delta u_{m}\|_{L^{2}}^{2}\right)-\int_{\Omega}\text{div}(a(x)\nabla u_{m}^{\prime})(-\Delta u_{m}^{\prime})dx+\int_{\Omega}u_{m}^{5}(-\Delta u_{m}^{\prime})dx=0.$
Estimating the Linear Terms and the Damping Commutator.
The term corresponding to the Kelvin-Voigt damping requires careful handling because the damping coefficient $a(x)$ is localized and non-constant. Expanding the divergence, we have:
$\int_{\Omega}\text{div}(a(x)\nabla u_{m}^{\prime})(-\Delta u_{m}^{\prime})dx=\int_{\Omega}a(x)|\Delta u_{m}^{\prime}|^{2}dx+\int_{\Omega}(\nabla a\cdot\nabla u_{m}^{\prime})\Delta u_{m}^{\prime}dx.$
The first term on the right-hand side provides the strong parabolic dissipation. The second term is a commutator of lower order. To absorb this commutator without losing derivatives, we strongly rely on the structural assumption on the damping function:
$|\nabla a(x)|^{2}\leq C_{a}a(x).$ (14)
Applying the Cauchy-Schwarz inequality and Young’s inequality taking (14) into account, we estimate:
$\left|\int_{\Omega}(\nabla a\cdot\nabla u_{m}^{\prime})\Delta u_{m}^{\prime}dx\right|\leq\int_{\Omega}\frac{|\nabla a|}{\sqrt{a}}|\nabla u_{m}^{\prime}|\left(\sqrt{a}|\Delta u_{m}^{\prime}|\right)dx$
$\leq\frac{1}{2}\int_{\Omega}a(x)|\Delta u_{m}^{\prime}|^{2}dx+\frac{C_{a}}{2}\int_{\Omega}|\nabla u_{m}^{\prime}|^{2}dx.$
Absorbing the dissipative term, the linear part of the equation yields the differential inequality:
$Y_{m}^{\prime}(t)+\frac{1}{2}\int_{\Omega}a(x)|\Delta u_{m}^{\prime}|^{2}dx\leq C_{a}Y_{m}(t)+\int_{\Omega}u_{m}^{5}\Delta u_{m}^{\prime}dx.$
Controlling the Quintic Nonlinearity via Gagliardo-Nirenberg
We now face the most delicate part of the proof: estimating the nonlinear source term uniformly. Integrating by parts once, we have:
$\int_{\Omega}u_{m}^{5}\Delta u_{m}^{\prime}dx=-5\int_{\Omega}u_{m}^{4}\nabla u_{m}\cdot\nabla u_{m}^{\prime}dx.$
To bound this term, we apply the generalized Hölder’s inequality with exponents $p_{1}=3$, $p_{2}=6$, and $p_{3}=2$:
$\left|5\int_{\Omega}u_{m}^{4}\nabla u_{m}\cdot\nabla u_{m}^{\prime}dx\right|\leq 5\|u_{m}^{4}\|_{L^{3}}\|\nabla u_{m}\|_{L^{6}}\|\nabla u_{m}^{\prime}\|_{L^{2}}=5\|u_{m}\|_{L^{12}}^{4}\|\nabla u_{m}\|_{L^{6}}\|\nabla u_{m}^{\prime}\|_{L^{2}}.$
Let us analyze each term on the right-hand side in terms of the strong energy $Y_{m}(t)$:Clearly, $\|\nabla u_{m}^{\prime}\|_{L^{2}}\leq\sqrt{2}Y_{m}(t)^{1/2}$.By the continuous Sobolev embedding $H^{2}(\Omega)\hookrightarrow W^{1,6}(\Omega)$ in dimension 3, we have
$\|\nabla u_{m}\|_{L^{6}}\leq C\|\Delta u_{m}\|_{L^{2}}\leq CY_{m}(t)^{1/2}$
The core difficulty lies in controlling $\|u_{m}\|_{L^{12}}^{4}$ efficiently to ensure a low polynomial growth for the resulting ODE. Instead of using the brute-force embedding $H^{2}\hookrightarrow L^{\infty}$ (which would yield a very high power and severely restrict the time of existence), we use the Gagliardo-Nirenberg interpolation inequality in $3D$. We interpolate $L^{12}$ between $L^{6}$ (which is uniformly bounded by the weak energy $E_{0}$) and $H^{2}$:
$\|u_{m}\|_{L^{12}}\leq C\|\Delta u_{m}\|_{L^{2}}^{1/4}\|u_{m}\|_{L^{6}}^{3/4}.$
Raising this inequality to the 4th power, and recalling that $\|u_{m}\|_{L^{6}}\leq C\|\nabla u_{m}\|_{L^{2}}\leq CE_{0}^{1/2}$, we obtain:
$\|u_{m}\|_{L^{12}}^{4}\leq C\|\Delta u_{m}\|_{L^{2}}\|u_{m}\|_{L^{6}}^{3}\leq CY_{m}(t)^{1/2}E_{0}^{3/2}.$
Combining all these estimates, the nonlinear term is bounded by:
$\left|\int_{\Omega}u_{m}^{5}\Delta u_{m}^{\prime}dx\right|\leq C\left(Y_{m}(t)^{1/2}E_{0}^{3/2}\right)\left(Y_{m}(t)^{1/2}\right)\left(Y_{m}(t)^{1/2}\right)=C_{1}E_{0}^{3/2}Y_{m}(t)^{3/2},$
where $C_{1}>0$ is a constant depending only on $\Omega$ and the Sobolev embeddings.
The Bernoulli Differential Inequality and Local Existence.
Putting the linear and nonlinear estimates together, we arrive at the following differential inequality for the strong energy:
$Y_{m}^{\prime}(t)\leq C_{a}Y_{m}(t)+C_{1}E_{0}^{3/2}Y_{m}(t)^{3/2}.$
This is a Bernoulli-type differential inequality. To solve it explicitly, we set $z(t)=Y_{m}(t)^{-1/2}$, which transforms the inequality into the linear form $z^{\prime}(t)\geq-\frac{C_{a}}{2}z(t)-\frac{C_{1}E_{0}^{3/2}}{2}$. Multiplying by the integrating factor $e^{C_{a}t/2}$ and integrating from 0 to $t$, we can trace back to $Y_{m}(t)$ to obtain:
$Y_{m}(t)\leq\frac{Y_{m}(0)e^{C_{a}t}}{\left[1-\frac{C_{1}E_{0}^{3/2}}{C_{a}}\sqrt{Y_{m}(0)}\left(e^{C_{a}t/2}-1\right)\right]^{2}}.$
Since $Y_{m}(0)$ is bounded by the $H^{2}\times H^{1}$ norm of the initial data, the denominator is strictly positive and bounded away from zero on some maximal interval $0,T_{max})$, where $T_{max}>0$ depends continuously on $Y_{m}(0)$, $E_{0}$, and the structural constant of the damping $C_{a}$. This uniform bound on $Y_{m}(t)$ over $[0,T]$ guarantees that the approximate solutions $u_{m}$ remain bounded in $L^{\infty}(0,T;H^{2}\cap H_{0}^{1})$ and $u_{m}^{\prime}$ in $L^{\infty}(0,T;H_{0}^{1})$. By standard compactness arguments (e.g., Aubin-Lions-Simon lemma), we can extract a subsequence that converges strongly, thus passing to the limit as $m\to\infty$ and proving the local well-posedness of strong solutions.
Theorem 3.1 (Local Well-posedness of Strong Solutions).
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with smooth boundary. Assume the damping coefficient $a\in W^{1,\infty}(\Omega)$ is non-negative and satisfies the structural condition $|\nabla a(x)|^{2}\leq C_{a}a(x)$ for almost every $x\in\Omega$.
For any initial data $(u_{0},u_{1})\in\left(H^{2}(\Omega)\cap H_{0}^{1}(\Omega)\right)\times H_{0}^{1}(\Omega)$, there exists a maximal time $T_{max}>0$, depending continuously on the initial weak energy $E_{0}$, the initial strong energy $Y(0)$, and the structural constant $C_{a}$, such that the critical wave equation admits a unique local strong solution $u$ on $[0,T]$ for any $T<T_{max}$.
Furthermore, the solution belongs to the strong regularity class:
$u$ $\in L^{\infty}(0,T;H^{2}(\Omega)\cap H_{0}^{1}(\Omega)),$
$u_{t}$ $\in L^{\infty}(0,T;H_{0}^{1}(\Omega)),$
$\sqrt{a(x)}\Delta u_{t}$ $\in L^{2}(0,T;L^{2}(\Omega)),$
and its strong energy $Y_{u}(t):=\frac{1}{2}\|\nabla u_{t}(t)\|_{L^{2}}^{2}+\frac{1}{2}\|\Delta u(t)\|_{L^{2}}^{2}$ satisfies the uniform bound:
$Y_{u}(t)\leq\frac{Y_{u}(0)e^{C_{a}t}}{\left[1-\frac{C_{1}E_{0}^{3/2}}{C_{a}}\sqrt{Y_{u}(0)}\left(e^{C_{a}t/2}-1\right)\right]^{2}},\quad\forall t\in[0,T_{max}).$ (15)
4. Uniqueness and Weak-Strong Uniqueness Principles
In this section, we address the uniqueness of the solutions obtained previously. It is a well-known open problem whether weak solutions in the basic energy space $C([0,T];H_{0}^{1}(\Omega))\cap C^{1}([0,T];L^{2}(\Omega))$ are unconditionally unique for the critical quintic wave equation in dimension 3. Because our Kelvin-Voigt damping $a(x)$ is localized, the dynamics in the undamped region $\{x\in\Omega:a(x)=0\}$ correspond exactly to the pure critical wave equation.
Therefore, without global Strichartz estimates, unconditional uniqueness is out of reach. Nevertheless, we can establish rigorous uniqueness within the class of strong solutions, as well as a Weak-Strong uniqueness principle relying on the auxiliary space $L^{4}(0,T;L^{12}(\Omega))$.
4.1. Uniqueness of Strong Solutions
We begin by showing that strong solutions, constructed via the Galerkin method for initial data in $H^{2}(\Omega)$, are uniquely determined.
Theorem 4.1 (Uniqueness of Strong Solutions).
Let $(u_{0},u_{1})\in(H^{2}(\Omega)\cap H_{0}^{1}(\Omega))\times H_{0}^{1}(\Omega)$. Suppose $u$ and $v$ are two strong solutions to the problem (1) in the class $L^{\infty}(0,T;H^{2}(\Omega)\cap H_{0}^{1}(\Omega))\cap W^{1,\infty}(0,T;H_{0}^{1}(\Omega))$ satisfying the same initial conditions. Then $u\equiv v$ on $[0,T]$.
Proof.
Let w=uvw=u-v. Then $w$ satisfies the initial value problem:
$w_{tt}-\Delta w-\text{div}(a(x)\nabla w_{t})+u^{5}-v^{5}=0,\quad\text{with}\quad w(0)=0,\quad w_{t}(0)=0.$ (16)
Testing the equation with $w_{t}$ (which is a valid operation since $w_{t}\in H_{0}^{1}(\Omega)$), and integrating over $\Omega$, we obtain the energy identity for the difference:
$\frac{1}{2}\frac{d}{dt}\left(\|w_{t}\|_{L^{2}}^{2}+\|\nabla w\|_{L^{2}}^{2}\right)+\int_{\Omega}a(x)|\nabla w_{t}|^{2}dx=-\int_{\Omega}(u^{5}-v^{5})w_{t}dx.$ (17)
By the Mean Value Theorem, we can bound the nonlinear difference as follows:
$|u^{5}-v^{5}|\leq 5\left(|u|^{4}+|v|^{4}\right)|w|.$
Since $u$ and $v$ are strong solutions in dimension 3, the Sobolev embedding $H^{2}(\Omega)\hookrightarrow L^{\infty}(\Omega)$ ensures that $u,v\in L^{\infty}(0,T;L^{\infty}(\Omega))$. Therefore, there exists a constant $M>0$, depending on the $L^{\infty}_{t}H^{2}_{x}$ norms of $u$ and $v$, such that:
$|u(t,x)|^{4}+|v(t,x)|^{4}\leq M\quad\text{for a.e. }(t,x)\in(0,T)\times\Omega.$
Substituting this back into the energy identity and dropping the non-negative dissipation term $\int a|\nabla w_{t}|^{2}dx$, we estimate the right-hand side using the Cauchy-Schwarz inequality:
$\frac{1}{2}\frac{d}{dt}\left(\|w_{t}\|_{L^{2}}^{2}+\|\nabla w\|_{L^{2}}^{2}\right)$ $\leq 5M\int_{\Omega}|w||w_{t}|dx$
$\leq\frac{5M}{2}\left(\|w\|_{L^{2}}^{2}+\|w_{t}\|_{L^{2}}^{2}\right).$
By the Poincaré inequality, $\|w\|_{L^{2}}^{2}\leq C_{P}\|\nabla w\|_{L^{2}}^{2}$. Thus, letting $E_{w}(t)=\frac{1}{2}\|w_{t}\|_{L^{2}}^{2}+\frac{1}{2}\|\nabla w\|_{L^{2}}^{2}$, we arrive at the linear differential inequality:
$E_{w}^{\prime}(t)\leq CE_{w}(t),$
where $C=5M\max(1,C_{P})$. Since $w(0)=w_{t}(0)=0$, we have $E_{w}(0)=0$. By Grönwall’s Lemma, $E_{w}(t)\equiv 0$ for all $t\in[0,T]$, which concludes the proof that $u\equiv v$. ∎
4.2. Uniqueness in the Strichartz Class
The true challenge arises when we relax the regularity of the solutions to the energy space. While unconditional uniqueness in the weak energy space $C([0,T];H_{0}^{1}(\Omega))\cap C^{1}([0,T];L^{2}(\Omega))$ remains a notorious open problem for the critical quintic wave equation in 3D, we can establish rigorous uniqueness within the auxiliary Strichartz space $L^{4}(0,T;L^{12}(\Omega))$. This is precisely the regularity class recovered by our well-posedness framework using the Littlewood-Paley frequency decomposition.
Theorem 4.2 (Uniqueness in the Strichartz Class).
Let $(u_{0},u_{1})\in H_{0}^{1}(\Omega)\times L^{2}(\Omega)$. Suppose $u$ and $v$ are two solutions to the critical problem (1) in the energy space $C([0,T];H_{0}^{1}(\Omega))\cap C^{1}([0,T];L^{2}(\Omega))$ sharing the same initial data. If, additionally, both solutions belong to the Strichartz class $u,v\in L^{4}(0,T;L^{12}(\Omega))$, then $u\equiv v$ on $[0,T]$.
Proof.
Let w=uvw=u-v. Since $u$ and $v$ share the same initial data, $w(0)=0$ and $w_{t}(0)=0$. Because both $u$ and $v$ belong to $L^{4}(0,T;L^{12}(\Omega))$, we can show that the nonlinear difference $u^{5}-v^{5}$ is regular enough to be tested against $w_{t}\in L^{\infty}(0,T;L^{2}(\Omega))$.
Using the algebraic inequality $|u^{5}-v^{5}|\leq C(|u|^{4}+|v|^{4})|w|$, we apply the generalized Hölder’s inequality in space with exact exponents $p_{1}=3$ and $p_{2}=6$:
$\|u^{5}(t)-v^{5}(t)\|_{L^{2}(\Omega)}$ $\leq C\left\|(|u(t)|^{4}+|v(t)|^{4})|w(t)|\right\|_{L^{2}(\Omega)}$
$\leq C\left(\|u(t)\|_{L^{12}(\Omega)}^{4}+\|v(t)\|_{L^{12}(\Omega)}^{4}\right)\|w(t)\|_{L^{6}(\Omega)}.$
By the Sobolev embedding $H_{0}^{1}(\Omega)\hookrightarrow L^{6}(\Omega)$, we have $\|w(t)\|_{L^{6}(\Omega)}\leq C_{S}\|\nabla w(t)\|_{L^{2}(\Omega)}$. Since $u,v\in L^{4}(0,T;L^{12}(\Omega))$, the function $h(t):=CC_{S}\left(\|u(t)\|_{L^{12}}^{4}+\|v(t)\|_{L^{12}}^{4}\right)$ belongs to $L^{1}(0,T)$. Thus, $u^{5}-v^{5}\in L^{1}(0,T;L^{2}(\Omega))$, which rigorously justifies taking the $L^{2}$-inner product of the weak equation for $w$ with $w_{t}$. This yields the energy identity:
$\frac{1}{2}\frac{d}{dt}\left(\|w_{t}\|_{L^{2}}^{2}+\|\nabla w\|_{L^{2}}^{2}\right)+\int_{\Omega}a(x)|\nabla w_{t}|^{2}dx=-\int_{\Omega}(u^{5}-v^{5})w_{t}dx.$ (18)
Dropping the non-negative dissipation term and defining the energy of the difference as $E_{w}(t)=\frac{1}{2}\|w_{t}\|_{L^{2}}^{2}+\frac{1}{2}\|\nabla w\|_{L^{2}}^{2}$, we estimate the right-hand side using the Cauchy-Schwarz inequality:
$E_{w}^{\prime}(t)$ $\leq\|u^{5}-v^{5}\|_{L^{2}(\Omega)}\|w_{t}\|_{L^{2}(\Omega)}$
$\leq h(t)\|\nabla w\|_{L^{2}(\Omega)}\|w_{t}\|_{L^{2}(\Omega)}$
$\leq h(t)E_{w}(t).$
Since $h\in L^{1}(0,T)$ and $E_{w}(0)=0$, Grönwall’s Lemma implies $E_{w}(t)\equiv 0$ for all $t\in[0,T]$. Consequently, $w\equiv 0$, meaning $u\equiv v$. ∎
5. Littlewood-Paley Strategy
In the present section the Littlewood–Paley decomposition reveals an intrinsic stabilizing mechanism: the Kelvin–Voigt operator behaves as a spectral smoothing device. Low frequencies absorb derivative losses via Bernstein inequalities, while high frequencies remain dispersive
5.1. The Philosophy: Frequencies Instead of Space
The central challenge is the Kelvin-Voigt term: $\text{div}(a(x)\nabla u_{t})$. By having two spatial derivatives ($\nabla\cdot\nabla$), this term belongs to the dual space $H^{-1}$. The classical Strichartz estimate for the wave equation does not accept terms in $H^{-1}$ on the right-hand side.
The spatial flaw: Attempting to isolate the damping with a spatial cutoff function $\psi(x)$ generates geometric commutators and creates a “no man’s land” ($0<\psi<1$) where the quintic nonlinearity loses control of the Strichartz norm.
The spectral solution: Instead of slicing the physical domain, we cut the frequency of the wave. The low-frequency spectral projection operator $P_{\leq N}$ possesses the property of Bernstein’s Inequality. It limits the cost of the derivatives, lifting the term from $H^{-1}$ back to $L^{2}$, a space where the Strichartz Theorem works perfectly.
5.2. Step 1: The Regularized Problem and the Energy
To apply the bootstrap argument (a priori estimate) without the nonlinearity blowing up, we truncate the quintic function. Let $f_{k}(s)$ be a globally Lipschitz function such that $f_{k}(s)=s^{5}$ for $|s|\leq k$.
We consider the strong solution $u$ (the index $k$ is omitted) of the problem:
$u_{tt}-\Delta u-\text{div}(a(x)\nabla u_{t})+f_{k}(u)=0$ $\text{in }\Omega\times(0,T),$ (19)
$u=0$ $\text{on }\partial\Omega\times(0,T),$ (20)
$u(0,x)=u_{0}(x),\quad u_{t}(0,x)=u_{1}(x)$ $\text{in }\Omega.$ (21)
Multiplying the PDE by $u_{t}$ and integrating over space and time from 0 to $t$, we obtain the Global Energy Identity:
$E(t)+\int_{0}^{t}\int_{\Omega}a(x)|\nabla u_{t}(s,x)|^{2}\,dx\,ds=E(0)\leq C_{0},$ (22)
where $E(0)$ depends exclusively on the initial data $(u_{0},u_{1})$. This guarantees uniform bounds in $k$ for $u\in L^{\infty}(0,T;H_{0}^{1}(\Omega))$, $u_{t}\in L^{\infty}(0,T;L^{2}(\Omega))$, and for the thermo-viscous dissipation.
5.3. Step 2: The Spectral Filter ($P_{\leq N}$)
We use the Spectral Theorem for the $\Delta$ operator with Dirichlet boundary conditions, see the Appendix of the present article. Let $\{\lambda_{j}\}$ be the eigenvalues and $\{\phi_{j}\}$ be the orthonormal eigenfunctions. For a threshold $N>0$, we define the Low-Frequency Projector:
$P_{\leq N}u=\sum_{\sqrt{\lambda_{j}}\leq N}(u,\phi_{j})_{L^{2}}\phi_{j}$
This operator commutes with $\Delta$ and $\partial_{t}$, and satisfies Bernstein’s Inequality:
$\|\nabla(P_{\leq N}g)\|_{L^{2}(\Omega)}\leq N\|g\|_{L^{2}(\Omega)}$ (23)
5.4. Step 3: Low Frequencies and Strichartz
We define the macroscopic part of the wave $w_{N}=P_{\leq N}u$. Applying $P_{\leq N}$ to the PDE (19):
$(w_{N})_{tt}-\Delta w_{N}=-P_{\leq N}(f_{k}(u))+P_{\leq N}\big(\text{div}(a(x)\nabla u_{t})\big)$ (24)
To tame the Kelvin-Voigt (KV) intrusive term, we transfer the derivative using the adjoint property and Bernstein’s inequality:
$\|P_{\leq N}\text{div}(a(x)\nabla u_{t})\|_{L^{2}}$ $=\sup_{\|\varphi\|_{L^{2}}=1}\langle a(x)\nabla u_{t},\nabla(P_{\leq N}\varphi)\rangle$
$\leq\sup_{\|\varphi\|_{L^{2}}=1}\|a(x)\nabla u_{t}\|_{L^{2}}\cdot(N\|\varphi\|_{L^{2}})$
$=N\|a(x)\nabla u_{t}\|_{L^{2}}$
To justify the interaction between the Kelvin–Voigt operator and the spectral projector, we exploit the selfadjointness of $P_{\leq N}$. For any $\phi\in L^{2}(\Omega)$, we write
$\big\langle P_{\leq N}\operatorname{div}(a\nabla u_{t}),\phi\big\rangle=\big\langle\operatorname{div}(a\nabla u_{t}),P_{\leq N}\phi\big\rangle.$ (25)
Integrating by parts in space yields
$\big\langle\operatorname{div}(a\nabla u_{t}),P_{\leq N}\phi\big\rangle=-\big\langle a\nabla u_{t},\nabla(P_{\leq N}\phi)\big\rangle.$ (26)
Using the Bernstein inequality for spectral projectors,
$\|\nabla(P_{\leq N}\phi)\|_{L^{2}(\Omega)}\leq CN\|\phi\|_{L^{2}(\Omega)},$ (27)
we obtain
$\|P_{\leq N}\operatorname{div}(a\nabla u_{t})\|_{L^{2}(\Omega)}\leq CN\|a\nabla u_{t}\|_{L^{2}(\Omega)}.$ (28)
Integrating in $L^{1}(0,T)$ and using the Global Energy (22):
$\|F_{KV}\|_{L^{1}L^{2}}\leq C_{a}N\sqrt{T}E_{0}^{1/2}$
With the right-hand side in $L^{1}L^{2}$, we apply Strichartz to $w_{N}$:
$\|w_{N}\|_{L^{5}L^{10}}\leq C_{S}E_{0}^{1/2}+C_{S}C_{a}N\sqrt{T}E_{0}^{1/2}+C_{1}\|u\|_{L^{5}L^{10}}^{5}$ (29)
where we used the standard Bernstein inequality for spectral projectors associated with the Dirichlet Laplacian (see Blair–Smith–Sogge).
5.5. Step 4: High Frequencies (The Microscopic Fluctuation)
Let $v_{N}=P_{>N}u=u-w_{N}$. To tame the quintic term, we need $\|v_{N}\|_{L^{5}L^{10}}$ to be small. Since the Sobolev embedding $H^{1}\hookrightarrow L^{10}$ fails in 3D, we exploit dispersion via initial data. Because the initial data $(u_{0},u_{1})$ does not depend on the truncation $k$, the projection onto high frequencies goes to zero as $N\to\infty$. Applying the classical Strichartz estimate to the high frequency, we guarantee that for short times:
$\|v_{N}\|_{L^{5}L^{10}}\leq\epsilon(N)$ (30)
where $\epsilon(N)\to 0$ as $N\to\infty$, independently of $k$.
5.6. Step 5: The Final Bootstrap Argument
To close the estimates and ensure that the nonlinearity does not blow up, we analyze the critical Strichartz norm of the solution. Let us denote this norm by $X(t):=\|u\|_{L^{5}(0,t;L^{10}(\Omega))}$.
By the triangle inequality, the norm of the full solution is bounded by the sum of its low-frequency and high-frequency components:
$X(T)\leq\|w_{N}\|_{L^{5}(0,T;L^{10}(\Omega))}+\|v_{N}\|_{L^{5}(0,T;L^{10}(\Omega))}.$ (31)
Substituting the explicit estimates (29) and (30) into (31), we gather all the components to obtain:
$X(T)\leq\underbrace{C_{S}E_{0}^{1/2}}_{\text{Initial data}}+\underbrace{\epsilon(N)}_{\text{High-freq. tail}}+\underbrace{C_{S}C_{a}N\sqrt{T}E_{0}^{1/2}}_{\text{Kelvin-Voigt penalty}}+\underbrace{C_{1}X(T)^{5}}_{\text{Nonlinearity}}.$ (32)
Here, the success of the argument hinges on a strict hierarchical choice of the parameters $N$ (the frequency threshold) and $T$ (the time horizon). We must proceed in the following precise order to avoid circular dependencies:
  1. (1)
    Fixing the Frequency $N$: The term $\epsilon(N)$ represents the energy of the high-frequency tail of the initial data. Since the initial data has finite energy, $\epsilon(N)\to 0$ as $N\to\infty$. We choose and fix a sufficiently large $N_{0}$ such that $\epsilon(N_{0})$ is arbitrarily small.
  2. (2)
    Fixing the Time $T$: With $N_{0}$ now fixed as a constant, the Kelvin-Voigt penalty term $C_{S}C_{a}N_{0}\sqrt{T}E_{0}^{1/2}$ solely depends on $T$. We can then choose the existence time $T>0$ sufficiently small so that this entire term is absorbed.
Once $N$ and $T$ are appropriately chosen, the linear terms and the small penalties can be consolidated into a single constant $K>0$. The inequality (32) then reduces to the standard polynomial form:
$X(T)\leq K+C_{1}X(T)^{5}.$ (33)
To conclude that $X(T)$ is bounded, we invoke a standard bootstrap principle (or continuity argument). Consider the function $p(x)=x-C_{1}x^{5}$. For a sufficiently small constant $K$ (which can be ensured by shrinking the initial time interval), the equation $x-C_{1}x^{5}=K$ has two positive real roots, say $r_{1}<r_{2}$. This creates a "forbidden region" $(r_{1},r_{2})$ where the inequality $x-C_{1}x^{5}\leq K$ cannot be mathematically satisfied.
Since the mapping $t\mapsto X(t)=\|u\|_{L^{5}(0,t;L^{10}(\Omega))}$ is absolutely continuous in time, and it starts at $X(0)=0$ (which is strictly less than $r_{1}$), the value of $X(t)$ can never jump across the forbidden region $(r_{1},r_{2})$ to reach the larger values. Therefore, $X(T)$ must remain safely trapped in the bounded interval $0,r_{1}]$ for all times $t\in[0,T]$.
This establishes a uniform bound for the $L^{5}_{t}L^{10}_{x}$ norm that depends only on the initial data and the structural constants of the problem, but is completely independent of the truncation parameter $k$. This uniform bound is the ultimate key that allows us to pass to the limit as $k\to\infty$ and deduce the local existence of the strong solution.
5.7. Step 6: Passing to the Limit ($k\to\infty$)
Recall that the truncated nonlinearity $f_{k}:\mathbb{R}\to\mathbb{R}$ is defined by (2). In particular, $f_{k}$ is odd, $C^{1}$, and satisfies the growth bounds
$|f_{k}(s)|\leq C\,(|s|^{5}+k^{4}|s|),\qquad|f_{k}^{\prime}(s)|\leq C\,k^{4},\qquad s\in\mathbb{R},$ (34)
with a constant $C$ independent of $k$.
Let $u_{k}$ be the weak solutions to the truncated problems on $(0,T)$ constructed in the previous step. From the uniform energy estimate we have
$u_{k}\ \text{bounded in }L^{\infty}(0,T;H_{0}^{1}(\Omega)),\qquad(u_{k})_{t}\ \text{bounded in }L^{\infty}(0,T;L^{2}(\Omega)).$ (35)
With the uniform bound established in the previous step, we can now pass to the limit in the truncated problem. The bootstrap argument ensures the existence of a time $T>0$ and a constant $C>0$, both independent of $k$, such that the sequence of solutions $\{u_{k}\}$ to the truncated problems (19) satisfies:
$\|u_{k}\|_{L^{\infty}(0,T;H_{0}^{1})}+\|(u_{k})_{t}\|_{L^{\infty}(0,T;L^{2})}+\|u_{k}\|_{L^{5}(0,T;L^{10})}\leq C.$ (36)
In particular, by Sobolev embedding in dimension $3$,
$\sup_{k}\,\|u_{k}\|_{L^{\infty}(0,T;L^{6}(\Omega))}\leq C(E_{0}).$ (37)
By the Banach-Alaoglu Theorem, we can extract a subsequence, still denoted by $\{u_{k}\}$, and find a limit function $u$ such that, as $k\to\infty$:
$u_{k}$ $\rightharpoonup u\quad\text{weak-*}\text{ in }L^{\infty}(0,T;H_{0}^{1}(\Omega)),$ (38)
$(u_{k})_{t}$ $\rightharpoonup u_{t}\quad\text{weak-*}\text{ in }L^{\infty}(0,T;L^{2}(\Omega)).$ (39)
Since $\{u_{k}\}$ is bounded in $L^{\infty}(0,T;H_{0}^{1}(\Omega))$ and its time derivative $\{(u_{k})_{t}\}$ is bounded in $L^{\infty}(0,T;L^{2}(\Omega))$, the Aubin-Lions Compactness Lemma implies that $\{u_{k}\}$ converges strongly to $u$ in $L^{2}(0,T;L^{2}(\Omega))$. Up to extracting a further subsequence, we obtain point-wise convergence almost everywhere:
$u_{k}(t,x)\to u(t,x)\quad\text{a.e. in }(0,T)\times\Omega.$ (40)
This point-wise convergence, combined with the definition of the truncation $f_{k}$, yields $f_{k}(u_{k})\to u^{5}$ almost everywhere. To pass to the limit in the weak formulation, we must control the nonlinear term. From the critical Strichartz bound in (36), we know that $u_{k}$ is uniformly bounded in $L^{5}(0,T;L^{10}(\Omega))$. Since $|f_{k}(s)|\leq|s|^{5}$, it follows that the sequence $\{f_{k}(u_{k})\}$ is uniformly bounded in $L^{1}(0,T;L^{2}(\Omega))$.
Applying standard weak convergence results (such as Lions’ Lemma), the almost everywhere convergence and the uniform integrability bound guarantee that:
$f_{k}(u_{k})\rightharpoonup u^{5}\quad\text{weakly in }L^{1}(0,T;L^{2}(\Omega)).$ (41)
We now identify the limit of the nonlinear term. Set
$g_{k}:=f_{k}(u_{k}),\qquad g:=u^{5}.$
We prove that, up to a subsequence,
$f_{k}(u_{k})\rightharpoonup u^{5}\quad\text{weakly in }L^{\infty}(0,T;L^{6/5}(\Omega))\quad\text{(hence also weakly in }L^{1}(0,T;L^{2}(\Omega))\text{).}$ (42)
Lemma 5.1 (Uniform integrability and weak compactness).
The family $\{f_{k}(u_{k})\}$ is bounded in $L^{\infty}(0,T;L^{6/5}(\Omega))$.
Proof.
Using (34) and Hölder,
$\|f_{k}(u_{k}(t))\|_{L^{6/5}}\leq C\big(\|u_{k}(t)\|_{L^{6}}^{5}+k^{4}\|u_{k}(t)\|_{L^{6/5}}\big).$
Since $\Omega$ is bounded, $\|u_{k}\|_{L^{6/5}}\leq C_{\Omega}\|u_{k}\|_{L^{6}}$, and by (37) the first term is uniformly bounded. For the second term, note that $f_{k}(s)=s^{5}$ when $|s|\leq k$ and $|f_{k}(s)|\leq C|s|^{5}$ for all $s$ (indeed, the truncation does not increase the magnitude), hence
$\|f_{k}(u_{k}(t))\|_{L^{6/5}}\leq C\|u_{k}(t)\|_{L^{6}}^{5}\leq C(E_{0}),$
uniformly in $k$ and $t$. This yields the claim. ∎
Lemma 5.2 (Identification of the limit).
Up to a subsequence,
$f_{k}(u_{k})\rightharpoonup u^{5},$
weakly in $L^{\infty}(0,T;L^{6/5}(\Omega))$.
Proof.
We decompose
$f_{k}(u_{k})-u^{5}=\big(f_{k}(u_{k})-u_{k}^{5}\big)+\big(u_{k}^{5}-u^{5}\big).$ (43)
Step 1: $u_{k}^{5}\rightharpoonup u^{5}$. By (37), $\{u_{k}^{5}\}$ is bounded in $L^{\infty}(0,T;L^{6/5}(\Omega))$. Together with the a.e. convergence $u_{k}\to u$ a.e., we have $u_{k}^{5}\to u^{5}$ a.e. Hence, by weak compactness and the standard a.e. identification principle (Vitali/Dunford–Pettis, or Lions’ lemma), we obtain
$u_{k}^{5}\rightharpoonup u^{5}\quad\text{weakly in }L^{\infty}(0,T;L^{6/5}(\Omega)).$ (44)
Step 2: the truncation error vanishes. Since $f_{k}(s)=s^{5}$ for $|s|\leq k$, we have
$f_{k}(u_{k})-u_{k}^{5}=\big(f_{k}(u_{k})-u_{k}^{5}\big)\,\mathbf{1}_{\{|u_{k}|>k\}}.$
Moreover, by the explicit definition (2), there exists $C>0$ such that
$|f_{k}(s)-s^{5}|\leq C\,|s|^{5}\,\mathbf{1}_{\{|s|>k\}}\qquad\text{for all }s\in\mathbb{R}.$
Therefore,
$\|f_{k}(u_{k}(t))-u_{k}(t)^{5}\|_{L^{6/5}}^{6/5}\leq C\int_{\{|u_{k}(t)|>k\}}|u_{k}(t,x)|^{6}\,dx.$
Using (37) and Chebyshev’s inequality,
$\int_{\{|u_{k}(t)|>k\}}|u_{k}(t)|^{6}\,dx\leq\|u_{k}(t)\|_{L^{6}}^{6}\leq C(E_{0}),$
and moreover the sets $\{|u_{k}(t)|>k\}$ shrink as $k\to\infty$. Hence for a.e. $t$,
$\int_{\{|u_{k}(t)|>k\}}|u_{k}(t)|^{6}\,dx\longrightarrow 0,\quad\text{as }k\to\infty.$
By dominated convergence in time (using the uniform $L^{\infty}_{t}L^{6}_{x}$ bound), we deduce
$\|f_{k}(u_{k})-u_{k}^{5}\|_{L^{1}(0,T;L^{6/5}(\Omega))}\longrightarrow 0.$ (45)
Combining (43), (44), and (45), we conclude that $f_{k}(u_{k})\rightharpoonup u^{5}$ weakly in $L^{\infty}(0,T;L^{6/5}(\Omega))$, proving (42). ∎
Finally, for any test function $\varphi\in L^{1}(0,T;L^{6}(\Omega))$ we have
$\int_{0}^{T}\!\!\int_{\Omega}f_{k}(u_{k})\,\varphi\,dx\,dt\longrightarrow\int_{0}^{T}\!\!\int_{\Omega}u^{5}\,\varphi\,dx\,dt,$
which is the desired convergence in the weak formulation.
Finally, the linear terms pass to the limit straightforwardly via weak convergence. The thermo-viscous damping term converges weakly since $\sqrt{a(x)}\nabla(u_{k})_{t}$ is uniformly bounded in $L^{2}(0,T;L^{2}(\Omega))$. Thus, the limit function $u$ satisfies the original critical wave equation with Kelvin-Voigt damping in the sense of distributions on $(0,T)\times\Omega$. This completes the proof of local existence.
Theorem 5.3 (Local Well-Posedness for Arbitrary Data via Spectral Filtering).
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with smooth boundary, and $a\in W^{1,\infty}(\Omega)$ be a non-negative damping coefficient. For any arbitrarily large initial data $(u_{0},u_{1})\in H_{0}^{1}(\Omega)\times L^{2}(\Omega)$, there exists a time $T>0$ (depending on the frequency profile of the initial data) such that the critical Kelvin-Voigt damped wave equation (1) admits a unique local solution $u$ on $[0,T]$.
Furthermore, the solution satisfies the energy identity and belongs to the intersection of the fundamental energy space and the critical Strichartz classes:
$u$ $\in C([0,T];H_{0}^{1}(\Omega))\cap L^{5}(0,T;L^{10}(\Omega))\cap L^{4}(0,T;L^{12}(\Omega)),$
$u_{t}$ $\in C([0,T];L^{2}(\Omega)),$
$\sqrt{a(x)}\nabla u_{t}$ $\in L^{2}(0,T;L^{2}(\Omega)).$
Remark 5.4.
The inclusion of the solution in the Strichartz space $L^{4}(0,T;L^{12}(\Omega))$ is a direct consequence of the admissible pairs in the Strichartz estimates applied during the bootstrap argument. Crucially, this is exactly the regularity required to trigger the Uniqueness Principle established in Theorem 4.2, thus elevating the result from mere existence to full well-posedness for large data.
5.8. Step 2: Uniform Higher Regularity Estimates via Strichartz
The primary objective of this section is to establish the existence of strong solutions for the critical quintic wave equation with Kelvin-Voigt damping, provided the initial data is sufficiently regular. This higher regularity is the cornerstone that will allow us, via a standard density argument, to obtain weak solutions for finite energy data later on.
The main result of this section is summarized in the following theorem:
Theorem 5.5 (Global Strong Solutions for Regular Data).
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with smooth boundary. Assume the damping coefficient $a\in W^{1,\infty}(\Omega)$ satisfies $a(x)\geq 0$ and the geometric condition $|\nabla a|^{2}\leq ca(x)$ for some constant $c>0$. For any regular initial data $(u_{0},u_{1})\in(H^{2}(\Omega)\cap H_{0}^{1}(\Omega))\times H_{0}^{1}(\Omega)$, the critical wave equation with Kelvin-Voigt damping admits a unique global strong solution $u$ satisfying:
$u\in C([0,T];H^{2}(\Omega)\cap H_{0}^{1}(\Omega))\cap C^{1}([0,T];H_{0}^{1}(\Omega)),$ (46)
along with the critical Strichartz regularity $u\in L^{4}(0,T;L^{12}(\Omega))$. Furthermore, the higher-order energy obeys the uniform a priori bound:
$\sup_{t\in[0,T]}\left(\|\nabla\partial_{t}u(t)\|_{2}^{2}+\|\Delta u(t)\|_{2}^{2}\right)\leq\mathcal{M}(E_{2}(0),E(0)),$ (47)
where $E_{2}(0)=\|\nabla u_{1}\|_{2}^{2}+\|\Delta u_{0}\|_{2}^{2}$, $E(0)$ is the standard first-order energy, and $\mathcal{M}$ is a constant strictly independent of any truncation parameter.
Proof of Theorem 5.5.
To prove this theorem, we consider the regularized problem with a truncated nonlinearity $f_{k}(s)$. For any fixed truncation parameter $k>0$, the nonlinearity $f_{k}(s)$ is globally Lipschitz. Standard theory for semilinear wave equations ensures that, for smooth initial data $(u_{0},u_{1})\in(H^{2}(\Omega)\cap H_{0}^{1}(\Omega))\times H_{0}^{1}(\Omega)$, the regularized problem admits a unique strong solution $u_{k}\in C([0,T];H^{2}(\Omega)\cap H_{0}^{1}(\Omega))\cap C^{1}([0,T];H_{0}^{1}(\Omega))$.
Our goal in this step is to establish an a priori bound for the higher-order energy of $u_{k}$ that is strictly independent of the truncation parameter $k$. To achieve this, we will couple the higher-order energy identity with the uniform Strichartz bounds.
Let us define the second-order energy functional:
$E_{2}(t):=\|\nabla\partial_{t}u_{k}(t)\|_{2}^{2}+\|\Delta u_{k}(t)\|_{2}^{2}.$ (48)
Taking the $L^{2}(\Omega)$-inner product of the exact PDE for $u_{k}$ with the multiplier $\Delta\partial_{t}u_{k}$, we obtain the following identity:
$\frac{1}{2}\frac{d}{dt}E_{2}(t)$ $+\int_{\Omega}a(x)|\Delta\partial_{t}u_{k}|^{2}\,dx$
$=-\int_{\Omega}(\nabla a\cdot\nabla\partial_{t}u_{k})\Delta\partial_{t}u_{k}\,dx+\int_{\Omega}f_{k}^{\prime}(u_{k})\nabla u_{k}\cdot\nabla\partial_{t}u_{k}\,dx.$ (49)
For the term involving the gradient of the damping coefficient, we employ the geometric assumption $|\nabla a|^{2}\leq ca(x)$. By writing $|\nabla a|=\frac{|\nabla a|}{\sqrt{a}}\sqrt{a}$ and applying Young’s inequality, we deduce:
$\left|\int_{\Omega}(\nabla a\cdot\nabla\partial_{t}u_{k})\Delta\partial_{t}u_{k}\,dx\right|$ $\leq\int_{\Omega}\frac{|\nabla a|}{\sqrt{a(x)}}|\nabla\partial_{t}u_{k}|\cdot\sqrt{a(x)}|\Delta\partial_{t}u_{k}|\,dx$
$\leq\frac{1}{2}\int_{\Omega}a(x)|\Delta\partial_{t}u_{k}|^{2}\,dx+\frac{c}{2}\|\nabla\partial_{t}u_{k}\|_{2}^{2}.$ (50)
The first term on the right-hand side is absorbed by the strong dissipation on the left-hand side of (5.8), while the second term is bounded by $\frac{c}{2}E_{2}(t)$.
The most delicate part is the uniform estimate for the nonlinear term. Since $|f_{k}^{\prime}(s)|\leq 5|s|^{4}$ universally for all $k$, we can estimate the integral using the generalized Hölder inequality with exponents $(\frac{1}{3},\frac{1}{6},\frac{1}{2})$:
$\left|\int_{\Omega}f_{k}^{\prime}(u_{k})\nabla u_{k}\cdot\nabla\partial_{t}u_{k}\,dx\right|$ $\leq 5\int_{\Omega}|u_{k}|^{4}|\nabla u_{k}||\nabla\partial_{t}u_{k}|\,dx$
$\leq 5\|u_{k}\|_{12}^{4}\|\nabla u_{k}\|_{6}\|\nabla\partial_{t}u_{k}\|_{2}.$ (51)
In dimension $n=3$, the Sobolev embedding $H^{2}(\Omega)\cap H_{0}^{1}(\Omega)\hookrightarrow W^{1,6}(\Omega)$ ensures that $\|\nabla u_{k}\|_{6}\leq C_{sob}\|\Delta u_{k}\|_{2}$. Inserting this into (5.8) and using Young’s inequality, we obtain:
$\left|\int_{\Omega}f_{k}^{\prime}(u_{k})\nabla u_{k}\cdot\nabla\partial_{t}u_{k}\,dx\right|$ $\leq 5C_{sob}\|u_{k}\|_{12}^{4}\|\Delta u_{k}\|_{2}\|\nabla\partial_{t}u_{k}\|_{2}$
$\leq C\|u_{k}(t)\|_{12}^{4}\left(\|\Delta u_{k}(t)\|_{2}^{2}+\|\nabla\partial_{t}u_{k}(t)\|_{2}^{2}\right)$
$=C\|u_{k}(t)\|_{12}^{4}E_{2}(t).$ (52)
Combining all the estimates back into (5.8), we arrive at the differential inequality:
$\frac{d}{dt}E_{2}(t)+\int_{\Omega}a(x)|\Delta\partial_{t}u_{k}|^{2}\,dx\leq C\left(1+\|u_{k}(t)\|_{12}^{4}\right)E_{2}(t).$ (53)
Applying Gronwall’s Lemma over the interval $[0,t]$ yields:
$E_{2}(t)\leq E_{2}(0)\exp\left(Ct+C\int_{0}^{t}\|u_{k}(s)\|_{12}^{4}\,ds\right).$ (54)
Notice that the integral inside the exponential is exactly the fourth power of the Strichartz norm $\|u_{k}\|_{L^{4}(0,t;L^{12}(\Omega))}^{4}$. As demonstrated in the subsequent sections, the Strichartz analysis guarantees that $\|u_{k}\|_{L^{4}L^{12}}\leq C_{S}$, where $C_{S}$ depends only on the initial energy $E(0)$ and is entirely independent of $k$.
Consequently, we obtain a uniform higher-order energy bound:
$\sup_{t\in[0,T]}\left(\|\nabla\partial_{t}u_{k}(t)\|_{2}^{2}+\|\Delta u_{k}(t)\|_{2}^{2}\right)\leq\mathcal{M}(E_{2}(0),E(0)),$ (55)
where $\mathcal{M}$ is a constant independent of $k$. This uniform bound is the cornerstone that allows us to extract a strongly convergent subsequence as $k\to\infty$, thereby proving the existence of strong solutions for the original critical quintic wave equation.
6. Stabilization Mechanism and Microlocal Structure
In this section we clarify the stabilization mechanism underlying the Kelvin–Voigt model and its interaction with the microlocal defect measure framework. The argument relies on the contradiction strategy introduced in Lemma 3.1 of [9], adapted here to the truncated system. Since the demonstration is exactly the same in [9], simply replacing the frictional dissipation term with the Kelvin-Voigt type dissipative term, we only give an idea of the proof by showing the behavior of the Kelvin-Voigt term.
6.1. Contradiction setting
To prove the uniform exponential stabilization, we argue by contradiction. Suppose that the uniform observability inequality fails. Then, there exists a sequence of solutions $\{u_{k}^{m}\}_{m\in\mathbb{N}}$ (keeping the truncation parameter $k$ fixed for the moment) such that their energies are not uniformly controlled by the dissipation. To analyze the regime where compactness is lost, we introduce the normalized sequence:
$v_{k}^{m}(t,x)=\frac{u_{k}^{m}(t,x)}{\alpha_{m}},\quad\text{where }\alpha_{m}=\sqrt{E_{u_{k}^{m}}(0)}.$ (56)
By construction, this new sequence satisfies $E_{v_{k}^{m}}(0)=1$ for all $m$. Consequently, $\{v_{k}^{m}\}_{m\in\mathbb{N}}$ is uniformly bounded in the energy space $L^{\infty}(0,T;H^{1}_{0}(\Omega))\cap W^{1,\infty}(0,T;L^{2}(\Omega))$ and converges weakly (up to a subsequence) to some limit function $v_{k}$.
While the Rellich-Kondrachov theorem provides strong convergence in lower-order spaces (such as $L^{2}_{t,x}$), the kinetic energy norm $\|\partial_{t}v_{k}^{m}\|_{L^{2}}^{2}$ lacks compactness. To rigorously quantify this defect of compactness—specifically, the concentration of high-frequency oscillations—we employ the framework of Microlocal Defect Measures introduced by P. Gérard [16] and L. Tartar [30].
By the structure theorem for bounded sequences in $L^{2}$, there exists a non-negative Radon measure $\mu$ defined on the cosphere bundle $S^{*}(\Omega\times(0,T))$ such that, for any classical pseudo-differential operator $A$ of order zero with principal symbol $\sigma_{0}(A)$, we have:
$\lim_{m\to\infty}\langle A(\partial_{t}v_{k}^{m}),\partial_{t}v_{k}^{m}\rangle_{L^{2}(\Omega\times(0,T))}=\langle\mu,\sigma_{0}(A)\rangle.$ (57)
In particular, choosing $A$ as the multiplication by a smooth test function $\phi(t,x)$, equation (57) reveals that $\mu$ characterizes the weak-$\star$ limit of the kinetic energy density:
$|\partial_{t}v_{k}^{m}|^{2}\rightharpoonup\int_{S^{*}_{t,x}}d\mu\quad\text{in the sense of measures.}$ (58)
Once that uniform stabilization fails, then there exist sequences of solutions $(v_{k}^{m})$ for the truncated problem and initial data normalized such that
$E_{v_{k}}^{m}(0)=1,\qquad\int_{0}^{T}\int_{\Omega}a(x)|\nabla v_{k,t}^{m}|^{2}\,dxdt\longrightarrow 0,\penalty 10000\ \hbox{as}\penalty 10000\ m\rightarrow\infty\penalty 10000\ (k\penalty 10000\ \hbox{fixed}).$ (59)
while
$E_{k}^{m}(t)\geq c_{0}>0,\qquad\forall t\in[0,T].$ (60)
The truncation introduces two parameters:
  • $k$ controlling the Galerkin/Littlewood–Paley truncation,
  • $m$ associated with the contradiction sequence.
Initially, the constants in Lemma 3.1 depend on $k$, but this dependence is removed through a second contradiction argument.
6.2. Loss of strong convergence due to Kelvin–Voigt damping
Passing to the limit $m\to\infty$, the key difficulty appears in the Kelvin–Voigt term. Unlike frictional damping, we no longer recover strong convergence of the residual in $L^{1}(0,T;L^{2}(\Omega))$.
More precisely, from the equation we obtain only
$\partial_{t}\Box v_{k}^{m}\longrightarrow 0\quad\text{in }L^{1}_{\mathrm{loc}}(0,T;H^{-2}(\Omega)),$ (61)
where $\Box=\partial_{t}^{2}-\Delta$ denotes the wave operator.
This loss of one derivative is intrinsic to the Kelvin–Voigt mechanism. Indeed, recalling the scaling argument
$u_{\lambda}(t,x)=u(\lambda t,\lambda x),$
we observe that
$\Delta u_{t}\mapsto\lambda^{3}\Delta u_{t},$
revealing the higher-order nature of the viscous term.
6.3. Microlocal consequence
Estimate (61) implies only that the associated microlocal defect measure $\mu$ satisfies
$\mathrm{supp}\,\mu\subset\mathrm{Char}(P),$ (62)
where $P=\partial_{t}^{2}-\Delta$ is the principal wave operator.
At this level of regularity we cannot yet conclude propagation along bicharacteristics. However, outside the damping region $\omega$, the Kelvin–Voigt term vanishes and the equation reduces to the free wave equation.
Therefore, on $\Omega\setminus\omega$, standard microlocal propagation applies, yielding
$\mathrm{supp}\,\mu\text{ is a union of bicharacteristics of }P,$ (63)
and, consequently the support of $\mu$ is transported along generalized bicharacteristics.
6.4. Vanishing of the defect measure in the damping region
From (59), we recover
$a(x)\nabla u_{k,t}^{m}\longrightarrow 0\quad\text{strongly in }L^{2},$
which implies
$\mu=0\quad\text{on }\omega.$ (64)
Combining (63) and (64), we deduce:
If every bicharacteristic of $P$ intersects $\omega$, then $\mu\equiv 0$.
This is precisely the Geometric Control Condition (GCC).
6.5. Role of the CGC / GCC and Non-Invasive Geometries
Under the standard Geometric Control Condition (GCC), uniform stabilization follows immediately since the defect measure $\mu$ is transported along the bicharacteristic flow and trivially annihilated upon entering the damping region $\omega$.
However, in many physical scenarios—particularly those involving inhomogeneous media, variable coefficients, or complex boundary configurations—the classical localized GCC fails due to the existence of trapped rays. These are generalized bicharacteristics that oscillate indefinitely without ever entering a standard localized control region, allowing high-frequency energy to concentrate and consequently preventing uniform exponential decay.
To overcome this fundamental geometric obstruction without trivially damping the entire domain, one can appeal to the highly ubiquitous, “non-invasive” damping geometries developed in the works of Cavalcanti et al. [6, 7, 8]. The core strategy in these papers relies on the intricate topological construction of a dissipative region $\omega$ that acts as a pervasive geometrical net.
Specifically, it is possible to design an open set $\omega$ satisfying two seemingly antagonistic properties:
  • Dynamical interception: The region $\omega$ is strategically distributed such that it intersects every single generalized bicharacteristic within a uniform finite time $T_{0}>0$.
  • Arbitrarily small volume: The Lebesgue measure of the damping region can be made arbitrarily small, i.e., $\operatorname{meas}(\omega)<\epsilon$ for any given $\epsilon>0$.
From the microlocal perspective, this construction reveals a profound property of the defect measure framework: it is completely insensitive to volumetric and metric considerations. The measure $\mu$ strictly encodes high-frequency concentration along the dynamical flow. As long as the damping set $\omega$ dynamically intercepts all invariant sets of the flow, the relation $\mu=0$ on $\omega$ will systematically propagate along the bicharacteristics, ultimately forcing $\mu\equiv 0$ everywhere in the domain.
Therefore, even when the underlying non-homogeneous metric strictly forbids stabilization via classical localized damping, the robust microlocal structure of the Kelvin-Voigt dissipation presented in this paper can be seamlessly coupled with these geometrically complex, arbitrarily small-measure damping sets to guarantee global exponential stabilization.
6.6. Removal of the truncation parameter $k$
The first contradiction argument yields estimates with constants depending on $k$. A second contradiction eliminates this dependence. (see [9])
Assuming failure of uniform bounds in $k$, we construct a renormalized sequence. Repeating the defect-measure argument shows again that $\mu\equiv 0$, producing the final contradiction. Indeed,
In previous works dealing with subcritical or Lipschitz non-linearities (e.g., [9]), the dependence on the truncation parameter $k$ is typically removed through a second contradiction argument, assuming $C_{k}\to\infty$ and applying the microlocal defect measure framework once again.
Here, however, we can bypass this second contradiction entirely by exploiting the sharp Unique Continuation Property (UCP) for the wave equation, taking full advantage of the critical Strichartz estimates obtained in Section 2.
Assume, by contradiction, that the uniform observability inequality fails for the original problem. By extracting a normalized contradiction sequence and passing to the weak limit as done previously, we obtain a limit function $v$ satisfying:
$\begin{cases}\Box v+\alpha^{4}v^{5}=0&\text{in }\Omega\times(0,T),\\ \partial_{t}v=0&\text{in }\omega\times(0,T).\end{cases}$ (65)
Differentiating the equation with respect to time, the function $w=\partial_{t}v$ solves the linearized equation:
$\Box w+V(x,t)w=0,\quad\text{where }V(x,t)=5\alpha^{4}v^{4}.$ (66)
This is the crucial step: thanks to the well-posedness framework established in Section 2, the limit solution $v$ inherits the critical Strichartz regularity $v\in L^{4}(0,T;L^{12}(\Omega))$. Consequently, the potential $V(x,t)$ belongs exactly to the critical space $L^{1}(0,T;L^{3}(\Omega))$.
According to the sharp unique continuation result by Duyckaerts, Zhang, and Zuazua [14] (Theorem 2.2), the observability over $\omega\times(0,T)$ for a wave equation with a potential in this exact $L^{1}_{t}L^{3}_{x}$ class implies that $w\equiv 0$ in the whole domain $\Omega\times(0,T)$.
This immediately yields $v\equiv 0$. Since the limit solution vanishes identically, the equipartition of energy forces the initial normalized energy to go to zero, which contradicts the normalization assumption ($E_{v^{m}}(0)=1$). This single contradiction proves the uniform observability inequality directly for the critical problem, eliminating any dependence on the truncation parameter $k$.
Theorem 6.1 (Global Exponential Stabilization).
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with smooth boundary. Assume that the localized Kelvin-Voigt damping coefficient $a\in W^{1,\infty}(\Omega)$ is non-negative and strictly positive on an open subset $\omega\subset\Omega$. Suppose further that $\omega$ satisfies the Geometric Control Condition (GCC), or is constructed as a non-invasive arbitrarily small-measure damping set that dynamically intercepts all generalized bicharacteristics of the wave operator.
Then, for any arbitrarily large initial data $(u_{0},u_{1})\in H_{0}^{1}(\Omega)\times L^{2}(\Omega)$, the unique global solution $u$ of the critical Kelvin-Voigt damped wave equation (1) stabilizes exponentially. That is, there exist uniform constants $C\geq 1$ and $\gamma>0$, independent of the initial data, such that the total energy of the system, defined by
$E_{u}(t)=\frac{1}{2}\|u_{t}(t)\|_{L^{2}(\Omega)}^{2}+\frac{1}{2}\|\nabla u(t)\|_{L^{2}(\Omega)}^{2}+\frac{1}{6}\|u(t)\|_{L^{6}(\Omega)}^{6},$ (67)
satisfies the decay estimate:
$E_{u}(t)\leq CE_{u}(0)e^{-\gamma t},\quad\forall t\geq 0.$ (68)
6.7. Conclusion
The Kelvin–Voigt mechanism alters the regularity of the residual, reducing convergence from $H^{-1}$ to $H^{-2}$. This prevents direct propagation everywhere but preserves the microlocal structure outside the damping region.
The stabilization mechanism therefore remains entirely governed by:
  • Microlocal propagation in $\Omega\setminus\omega$,
  • Vanishing of $\mu$ in $\omega$,
  • Geometric interception of bicharacteristics.
Remark 6.2 (Absence of dissipation on the illuminated boundary).
An important geometric configuration arises when the dissipative mechanism is absent on the illuminated portion of the boundary.
In such situations, generalized bicharacteristics may undergo multiple reflections without necessarily intersecting the damping region. The analysis must therefore incorporate the microlocal structure of the reflected flow.
The asymptotic behavior of high-frequency energy concentration for the Dirichlet problem was investigated by Gérard and Leichtnam [17]. Their results describe the ergodic and propagation properties of eigenfunctions, providing a rigorous framework for understanding how energy distributes under repeated reflections at the boundary.
From the stabilization viewpoint, the defect-measure argument remains valid provided the reflected bicharacteristic flow does not admit trapped trajectories avoiding the dissipative region. In particular, if the underlying billiard dynamics prevents the formation of invariant concentration sets, one still obtains $\mu\equiv 0$.
Therefore, the stabilization mechanism developed in this work is compatible with configurations where dissipation is missing on part of the illuminated boundary, under suitable dynamical conditions on the generalized bicharacteristic flow.
Appendix A Spectral Projectors and Bernstein Inequalities in Bounded Domains
In $\mathbb{R}^{n}$, the classical Littlewood-Paley decomposition relies heavily on the Fourier transform. For bounded domains $\Omega\subset\mathbb{R}^{3}$ with Dirichlet boundary conditions, the natural substitute is the spectral decomposition associated with the Laplace operator. In this appendix, we recall the definitions of the low and high-frequency spectral projectors and establish the corresponding Bernstein-type inequalities.
Let $A=-\Delta$ with domain $D(A)=H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$. Since $\Omega$ is bounded, $A$ is a strictly positive, self-adjoint operator with a compact inverse. By the Spectral Theorem, there exists a sequence of eigenvalues $\{\lambda_{j}\}_{j=1}^{\infty}$ such that
$0<\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{j}\to\infty\quad\text{as }j\to\infty,$ (69)
and a corresponding family of eigenfunctions $\{\phi_{j}\}_{j=1}^{\infty}$ which forms an orthonormal basis for $L^{2}(\Omega)$ and an orthogonal basis for $H^{1}_{0}(\Omega)$.
For any function $u\in L^{2}(\Omega)$, we have the unique expansion $u=\sum_{j=1}^{\infty}(u,\phi_{j})_{L^{2}}\phi_{j}$. The fractional Sobolev spaces can be characterized via the spectral coefficients. In particular, for $u\in H^{1}_{0}(\Omega)$:
$\|\nabla u\|_{L^{2}(\Omega)}^{2}=(Au,u)_{L^{2}(\Omega)}=\sum_{j=1}^{\infty}\lambda_{j}|(u,\phi_{j})_{L^{2}}|^{2}.$ (70)
A.1. Low-Frequency Projector and Bernstein’s Inequality
For a given frequency threshold $N>0$, we define the low-frequency projector $P_{\leq N}:L^{2}(\Omega)\to L^{2}(\Omega)$ by:
$P_{\leq N}u=\sum_{\sqrt{\lambda_{j}}\leq N}(u,\phi_{j})_{L^{2}}\phi_{j}.$ (71)
The operator $P_{\leq N}$ is an orthogonal projection in $L^{2}(\Omega)$, and hence it is self-adjoint. Moreover, it commutes with both $\Delta$ and $\partial_{t}$.
The key feature of the low-frequency projector is that it smooths the function, allowing us to control higher-order derivatives at the cost of the parameter $N$.
Lemma A.1 (Bernstein Inequality for Low Frequencies).
For any $u\in L^{2}(\Omega)$, $P_{\leq N}u\in H^{1}_{0}(\Omega)$ and we have:
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}\leq N\|u\|_{L^{2}(\Omega)}.$ (72)
Proof.
Using the spectral characterization (70) and the definition of $P_{\leq N}$, we obtain:
$\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}^{2}$ $=\sum_{\sqrt{\lambda_{j}}\leq N}\lambda_{j}|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq\sum_{\sqrt{\lambda_{j}}\leq N}N^{2}|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq N^{2}\sum_{j=1}^{\infty}|(u,\phi_{j})_{L^{2}}|^{2}=N^{2}\|u\|_{L^{2}(\Omega)}^{2}.$
Taking the square root yields the desired inequality. ∎
Remark A.2 (Orthogonality in the Energy Space and Regularity).
We implicitly assume that the boundary $\partial\Omega$ is sufficiently smooth (e.g., of class $C^{2}$) so that the domain of the Dirichlet Laplacian is fully characterized by $D(A)=H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$. Furthermore, because the eigenfunctions $\phi_{j}$ are orthogonal with respect to the inner product in both $L^{2}(\Omega)$ and $H^{1}_{0}(\Omega)$, the projectors $P_{\leq N}$ and $P_{>N}$ are orthogonal projections in the energy space. In particular, for any $u\in H^{1}_{0}(\Omega)$, we have the exact exact decoupling of the gradient norm:
$\|\nabla u\|_{L^{2}(\Omega)}^{2}=\|\nabla(P_{\leq N}u)\|_{L^{2}(\Omega)}^{2}+\|\nabla(P_{>N}u)\|_{L^{2}(\Omega)}^{2}.$ (73)
This perfect splitting guarantees that separating the macroscopic wave $w_{N}$ from the microscopic fluctuation $v_{N}$ does not artificially inflate the total energy bounds during the a priori estimates.
A.2. High-Frequency Projector and Poincaré-type Inequality
The high-frequency projector $P_{>N}:L^{2}(\Omega)\to L^{2}(\Omega)$ is defined naturally as the orthogonal complement:
$P_{>N}u=u-P_{\leq N}u=\sum_{\sqrt{\lambda_{j}}>N}(u,\phi_{j})_{L^{2}}\phi_{j}.$ (74)
While low frequencies control derivative losses, high frequencies provide a gain in integrability when bounded in the energy space, acting as an enhanced Poincaré inequality.
Lemma A.3 (Reverse Bernstein Inequality for High Frequencies).
For any $u\in H^{1}_{0}(\Omega)$, we have:
$\|P_{>N}u\|_{L^{2}(\Omega)}\leq\frac{1}{N}\|\nabla u\|_{L^{2}(\Omega)}.$ (75)
Proof.
By Plancherel’s theorem and the definition of $P_{>N}$, we write:
$\|P_{>N}u\|_{L^{2}(\Omega)}^{2}=\sum_{\sqrt{\lambda_{j}}>N}|(u,\phi_{j})_{L^{2}}|^{2}=\sum_{\sqrt{\lambda_{j}}>N}\frac{\lambda_{j}}{\lambda_{j}}|(u,\phi_{j})_{L^{2}}|^{2}.$
Since $\sqrt{\lambda_{j}}>N\implies\lambda_{j}>N^{2}$, we have $\frac{1}{\lambda_{j}}<\frac{1}{N^{2}}$ for all terms in the sum. Thus,
$\|P_{>N}u\|_{L^{2}(\Omega)}^{2}$ $\leq\frac{1}{N^{2}}\sum_{\sqrt{\lambda_{j}}>N}\lambda_{j}|(u,\phi_{j})_{L^{2}}|^{2}$
$\leq\frac{1}{N^{2}}\sum_{j=1}^{\infty}\lambda_{j}|(u,\phi_{j})_{L^{2}}|^{2}=\frac{1}{N^{2}}\|\nabla u\|_{L^{2}(\Omega)}^{2}.$
Taking the square root finishes the proof. ∎